OpenStudy (anonymous):

need to find the c-values of f(c)=0 on the interval [0,pi] when the beginning function is f(x)=cos(2x)

OpenStudy (anonymous):

so 2x=(pi/2) or 3*pi/2 . giving x= (pi)/4 and (pi)*3/4

OpenStudy (turingtest):

\[f(x)=\cos(2x)=0\] solve for x and exclude answers that are not in [0,pi]